OpenStudy (anonymous):

Under certain conditions argon (Ar) gas diffuses at a rate of 3.2 centimeters per second. Under the same conditions, an unknown gas diffuses at a rate of 4.5 centimeters per second. What is the approximate molar mass of the unknown gas? 10 g/mol 20 g/mol 28 g/mol 56 g/mol

OpenStudy (joannablackwelder):

\[\frac{ R _{2} }{ R _{1} }=\sqrt{\frac{ M _{1} }{ M _{2} }}\] where R is rate and M is molar mass

OpenStudy (anonymous):

So R2=4.5 R1=3.2 & M1=39.948 M2=X?

OpenStudy (joannablackwelder):

Yup :)

OpenStudy (joannablackwelder):

Welcome to Open Study! :)